Let X₁,X₂ ,...,Xₙ be a random sample from a N(μ,σ²) with σ² = 25 and sample
mean X̅ = 60 . Find a 95% confidence interval for μ.
1
Expert's answer
2019-03-27T14:36:59-0400

For a single mean from a normal distribution with known variance, a two-sided, 100(1 – α)% confidence interval is calculated by


"\bar{X}-z_{\alpha/2}*{\sigma \over \sqrt{n}}\le\mu\le\bar{X}+z_{\alpha/2}*{\sigma \over \sqrt{n}}"

For a 95% confidence interval for μ


"z_{\alpha/2}=z_{0.025}=1.96"

We have that


"\bar{X}=60, \sigma^2=25, n"

Then


"60-1.96*{\sqrt{25} \over \sqrt{n}}\le\mu\le60+1.96*{\sqrt{25} \over \sqrt{n}}"

"60-{9.8 \over \sqrt{n}}\le\mu\le60+{9.8 \over \sqrt{n}}"


"95\% CI\ [60-{9.8 \over \sqrt{n}}, 60+{9.8 \over \sqrt{n}}]"

If n=16


"60-1.96*{\sqrt{25} \over \sqrt{16}}\le\mu\le60+1.96*{\sqrt{25} \over \sqrt{16}}"


"57.55\le\mu\le62.45"

"95\% CI\ [57.55, 62.45]"